What is the wavelength of a photon with a frequency of 6.85 ✕ 1016 s-1?
1
Expert's answer
2019-10-03T05:19:50-0400

"\lambda = \frac{c}{\nu}"

where "\lambda" - wavelength of photon, "m" ,

"c" - speed of light, "c \approx 3 \cdot 10^{8} \; m \cdot s^{-1}",

"\nu" - frequency of photon, "s^{-1}".

"\lambda = \frac{3 \cdot 10^{8}}{6.85 \cdot 10^{16}} \approx 4.38 \cdot 10^{-9} \; m = 4.38 \; nm"


Answer: 4.38 nm.